# A Basic Evaluation of Neural Networks Trained with the Error Diffusion Learning Algorithm

### v2

@@ Abstract
Purpose
This paper presents a contemporary formulation of Kaneko’s Error Diffusion Learning Algorithm (EDLA) framework and evaluates its effectiveness through parity check, regression, and image classification tasks. The EDLA is a biologically inspired learning algorithm that is an alternative to the conventional backpropagation algorithm for training artificial neural networks.
Design/methodology/approach
EDLA employs a single global error signal diffused across neural networks comprising paired excitatory-inhibitory sublayers, eliminating traditional layer-wise error backpropagation. This study evaluates EDLA’s effectiveness using benchmark tasks—parity check, regression, and image classification—by systematically varying the neuron count, network depth, and learning rates to assess its performance comprehensively.
Findings
The experimental results demonstrate that EDLA achieves consistently high accuracy across multiple benchmarks, highlighting its effectiveness as a training algorithm for neural networks. The choice of learning rate, neuron count, and network depth significantly influences EDLA’s efficiency and convergence speed. Analysis of internal network representations reveals meaningful feature extraction capabilities comparable to networks trained via traditional backpropagation.
Originality/value
This study introduces EDLA, a biologically plausible alternative to traditional backpropagation previously underrecognized due to language barriers. By providing a modern and comprehensive formulation of EDLA, systematically evaluating its performance, and presenting empirical evidence of its effectiveness, this study significantly enhances the visibility and accessibility of EDLA, contributing substantially to biologically inspired neural network training methodologies.
@@ 1 Introduction
Artificial neural networks have substantially advanced numerous research domains, including image processing, signal processing, and natural language processing. The remarkable success of these networks largely stems from the backpropagation method proposed by Rumelhart et al. (Rumelhart et al., 1986), which has become a standard method for training neural networks across diverse architectures.
The theoretical foundation of neural networks can be traced back to the linear threshold unit (neuron) introduced by McCulloch and Pitts (Kleene, 1956; Mcculloch and Pitts, 1943). The fundamental mechanism of this neuron is based on the all-or-none law (Adrian, 1914; Kato, 1926) and still underlies modern neural network architectures. Rosenblatt’s development of the perceptron (Rosenblatt, 1958) represents another critical milestone in neural networks, marking the first realization of a trainable artificial neuron. However, a single-layer structure of perceptrons does not inherently have the capability to create nonlinear decision boundaries. Although multilayer perceptrons (MLPs) resolve this limitation, training these deeper networks has been a substantial challenge. The introduction of backpropagation has provided a solution for training MLPs. Subsequently, it has become the dominant paradigm for training various deep neural network architectures.
Despite the success of backpropagation, its biological plausibility has been questioned (Crick, 1989; Mostafa et al., 2018). This gap between artificial and biological neural systems has motivated the development of alternative learning algorithms. One example is Kaneko’s Error Diffusion Learning Algorithm (EDLA) (Kaneko, 1999). EDLA introduces a biologically inspired architecture that incorporates both excitatory and inhibitory synapses, reflecting biological neural systems more closely. The distinctive feature of this algorithm is its error-diffusion mechanism. In this approach, the difference between the network output and target values is diffused throughout the network during training, resulting in efficient training of sigmoid neural networks. Unlike standard backpropagation, the EDLA diffuses only a single error signal throughout the network, eliminating the need for backward propagation and error calculation for each layer. This simplification addresses one of the key problems highlighted by Mostafa (2018), which artificial neural networks based on backpropagation require buffering past presynaptic activities and synchronizing them with the corresponding error signals propagated sequentially from higher layers.
@@ 2 Error Diffusion Learning Algorithm (EDLA)
The Error Diffusion Learning Algorithm (EDLA) is a novel approach to neural network learning initially proposed by Kaneko (Kaneko, 1999). Although the original model is available only in Japanese, this section provides a comprehensive reinterpretation of EDLA.
@@ 2.2 Learning Rule and Mathematical Formulation
$z_{j}^{p,(1)}=g(a_{j}^{p,(1)})\quad\text{and}\quad z_{j}^{n,(1)}=g(a_{j}^{n,(1)})$ (3)
where $g(\cdot)$ is the activation function. In subsequent hidden layers $(l=2,...,L)$ where $L$ refers to the final hidden layer, the neuron outputs are given by
@@
The objective of training an EDLA network is to identify the optimal parameter set $\mathbf{W}$ that minimizes a predefined error function $E(\mathbf{W})$, similar to traditional standard MLPs. Gradient descent is a widely used optimization method in which parameters are iteratively updated in a direction that locally reduces the value of $E(\mathbf{W})$. In practice, most practitioners use stochastic gradient descent (SGD) (LeCun et al., 2015), which computes gradient estimates efficiently using subsets of training data (mini-batches). The EDLA also employs a method based on SGD.
@@
A key distinction between EDLA and traditional backpropagation is EDLA’s simplified approximation of error computations. Whereas traditional backpropagation individually computes and propagates neuron-specific errors $\delta$s from subsequent layers, EDLA diffuses only one global error signal $d$ uniformly across all layers. This diffusion mechanism significantly reduces complexity by eliminating the need for individual neuron-level error calculations, highlighting a unique feature of the EDLA approach compared with traditional backpropagation methods.
@@
where $\eta$ denotes the learning rate. Thus, the global error signal directly determines whether the positive or negative neuron weights are adjusted during the training.
@@
When the network output is smaller than the target value ($d>0$), the absolute values of the weights originating from positive neurons increase due to the properties $g^{\prime}(a)>0$ and $z>0$. Such weight adjustments amplify the outputs of positive neurons in subsequent layers while simultaneously suppressing the outputs of negative neurons in those layers. Consequently, these modifications reduce the discrepancy between the output and target value of the network. Conversely, when the network output exceeds the target value ($d<0$), the absolute values of the weights originating from negative neurons increase. These weight adjustments decrease the outputs of positive neurons in subsequent layers while simultaneously increasing the outputs of negative neurons. Thus, regardless of the sign of the global error signal, EDLA systematically reduces the difference between the network output and target values through successive weight updates. Importantly, this characteristic learning behavior of EDLA is maintained as long as $g^{\prime}(a)>0$, $z>0$, and the sign of the derivative of the loss function properly encodes whether the network output is below or above the target value.
For the EDLA, the activation function is typically chosen as the sigmoid function. It constrains the output of the neuron between 0 and 1 and its derivative between 0 and 0.25. By limiting both the neuron outputs and their derivatives, the sigmoid activation function provides stability to the training dynamics, avoiding huge weight updates and facilitating stable convergence.
@@ 2.3 Biological Interpretation
In the EDLA framework, the global error signal $d$ functions as a plasticity-guiding neuromodulatory signal, orchestrating synaptic weight updates through the coordinated activities of the presynaptic and postsynaptic neurons. This concept is strongly reminiscent of Hebbian learning (Hebb, 1949), which posits that changes in synaptic strength are driven by correlated patterns of neuronal activation. Specifically, the EDLA learning rule, represented by $\eta dg^{\prime}(a^{p,(l)}_{j})z^{p,(l-1)}_{i}$ for synaptic weights $w^{pp,(l)}_{ji}$, is proportional to the product of the derivative of the postsynaptic neuron’s activation function $g^{\prime}(a^{p,(l)}_{j})$ and the output of the presynaptic neuron $z^{p,(l-1)}_{i}$. This mathematical relationship can be interpreted as a Hebbian-like plasticity mechanism, wherein weight adjustments reflect a direct interaction between a presynaptic neuron (characterized by $z^{p,(l-1)}_{i}$ and a postsynaptic neuron (captured by $g^{\prime}(a^{p,(l)}_{j})$).
Moreover, EDLA employs a global error signal diffusion mechanism, in which the error signal $d$ uniformly propagates throughout the entire network. This diffusion process is conceptually analogous to biological neuromodulation and closely parallels the reward prediction error (RPE) mechanism observed in biological neural systems, which guides synaptic plasticity across neural circuits through diffuse neuromodulatory signals (Roelfsema and Holtmaat, 2018). This neuromodulatory-inspired global diffusion effectively orchestrates all synaptic modifications in the network; thereby, this mirrors biological processes wherein neuromodulators such as dopamine broadly regulate neural plasticity based on RPE signals.
@@ 2.4 Limitation
The inherent simplicity and approximation underlying EDLA introduce significant constraints, particularly concerning the network’s capability to handle multiple outputs effectively. In the EDLA framework, the global error signal $d$ is uniformly diffused throughout the network architecture. This global error diffusion inherently directs the optimization process toward minimizing the error associated with a single output node. Consequently, when dealing with single-output scenarios, the EDLA can efficiently and accurately specialize the network weights for a specific output. However, when an EDLA network contains multiple outputs, each output generates its own global error signal, which propagates throughout the network. Such simultaneous diffusion of distinct error signals leads to conflicting weight adjustments, as the network simultaneously attempts to minimize errors for multiple outputs. These conflicting updates inhibit the network’s ability to extract and generalize meaningful features from the input data, thereby undermining learning efficacy.
@@ 2.5 Extending EDLA to Multiple Outputs
The inherent single-output constraint of the EDLA represents a significant limitation when addressing tasks that require multiple outputs. A simple approach to adapting EDLA to overcome this limitation involves constructing multiple EDLA networks arranged in parallel. This parallelized architecture addresses a task that requires $K$ outputs by employing $K$-independent EDLA networks. Each individual network retains the fundamental EDLA architecture and operates autonomously. Input vector $\mathbf{x}$ is simultaneously provided to all networks, ensuring that they process the same input data. Each network independently computes output $y_{k}$, where $k\in\{1,\ldots,K\}$. In classification tasks, if the output $y_{k}$ is the maximum value among the outputs of the EDLA networks, $k$ is regarded as the overall output label.
@@
The weights of each individual EDLA network are independently optimized using the standard EDLA update rule. Consequently, this parallel architecture retains the capability, simplicity, and core principles inherent to single-output EDLA models, effectively extending EDLA’s capability to handle multioutput tasks. However, this parallelization approach increases computational costs because the total number of EDLA networks scales linearly with the required number of outputs. This trade-off shows the challenge of maintaining the computational simplicity of single-output EDLA networks while contending with the additional resource demands necessary for multioutput learning.
@@
Employing multiple single-output networks in an EDLA framework may seem computationally inefficient compared with traditional multi-output MLP architectures. Nevertheless, this parallelization method is reminiscent of module structures, which are organizational principles observed in biological neural systems, particularly in the brain (Casanova and Casanova, 2018). Mountcastle’s early neurophysiological research (Mountcastle, 1957) has demonstrated that neurons within the primate somatosensory cortex are organized into vertically aligned columns, each dedicated to a particular modality or receptive field. Additionally, neurons in the primary visual cortex are organized into modular structures known as hypercolumns, each specializing in processing specific features of visual stimuli such as orientation (Hubel and Wiesel, 1968). Each hypercolumn functions semi-independently, collectively contributing to overall visual perception by handling distinct components of the visual scene. Similarly, the inferior temporal cortex exhibits clusters of neurons that are selectively responsive to complex visual features including distinct object shapes (Sato et al., 2008; Tanaka, 1996). This clustered arrangement enables the efficient processing of diverse visual elements while simultaneously facilitating effective integration across features and categories. Another notable example is the barrel cortex of rodents, in which neural clusters (called barrels) are dedicated to independently processing tactile inputs from individual whiskers (Li and Crair, 2011; Woolsey and Van der Loos, 1970). Each barrel independently processes sensory information corresponding to its associated whisker, and the combined neural activity across the barrels generates a comprehensive tactile representation. These modular and distributed neural processing strategies enable the brain to handle complex and heterogeneous tasks efficiently while preserving both flexibility and robustness. Although the parallelized architecture of EDLA, which consists of multiple single-output networks dedicated to specific outputs, may be computationally less efficient than traditional multi-output MLPs, it reflects these biological modular architectures.

### v3

@@ Abstract
Background / introduction This paper presents a comprehensive formulation of Kaneko’s Error Diffusion Learning Algorithm (EDLA) and evaluates its effectiveness across parity check, regression, and image classification tasks. EDLA is a biologically inspired learning algorithm that provides an alternative to conventional backpropagation for training artificial neural networks.
Methods EDLA employs a single global error signal that diffuses across networks composed of paired positive and negative sublayers, eliminating traditional layer-wise error backpropagation. This study evaluates EDLA’s effectiveness using benchmark tasks, such as parity check, regression, and image classification, by systematically varying the neuron count, network depth, and learning rates to assess its performance comprehensively.
Results The experimental results demonstrate that EDLA achieves consistently high accuracy across multiple benchmarks, highlighting its effectiveness as a learning algorithm for neural networks. The choice of learning rate, neuron count, and network depth significantly influences EDLA’s efficiency and convergence speed. Analysis of internal network representations reveals meaningful feature extraction capabilities, and the network’s overall performance is found to be competitive with networks trained via conventional backpropagation, especially in shallow architectures.
Conclusions This study introduces EDLA, a biologically plausible alternative to traditional backpropagation previously underrecognized due to language barriers. By reformulating EDLA, systematically evaluating its performance, and presenting empirical evidence of its effectiveness, this study increases the visibility and accessibility of EDLA and contributes to biologically inspired training methodologies.
keywords:
Neural network, Learning algorithm, Error diffusion
@@ 1 Introduction
Artificial neural networks have substantially advanced numerous research domains, including image processing, signal processing, and natural language processing. The remarkable success of these networks largely stems from the backpropagation method proposed by Rumelhart et al. [1], which has become a standard method for training neural networks across diverse architectures.
The theoretical foundation of neural networks traces back to the linear threshold unit (neuron) introduced by McCulloch and Pitts [2, 3]. The fundamental mechanism of this neuron is based on the all-or-none law [4, 5] and still underlies modern neural network architectures. Rosenblatt’s development of the perceptron [6] represents another critical milestone in neural networks, marking the first realization of a trainable artificial neuron. However, because this trainable perceptron consists of only a single computational layer, it cannot inherently form nonlinear decision boundaries. Although multilayer perceptrons (MLPs) resolve this limitation, training these deeper networks has been a substantial challenge. The introduction of backpropagation has provided a solution for training MLPs. Subsequently, it has become the dominant paradigm for training various deep neural network architectures.
Despite the success of backpropagation, questions remain about its biological plausibility [7, 8]. This gap between artificial and biological neural systems has driven the development of alternative learning algorithms. One example is Kaneko’s Error Diffusion Learning Algorithm (EDLA) [9]. EDLA introduces a biologically inspired architecture that incorporates both excitatory and inhibitory synapses, reflecting biological neural systems more closely. A distinctive feature of this algorithm is its error-diffusion mechanism. In this approach, the difference between the network output and target values is diffused throughout the network during training, and has been reported to enable effective training with sigmoid activations. Unlike standard backpropagation, EDLA diffuses only a single error signal throughout the network, eliminating the need for backward propagation and error calculation for each layer. This simplification addresses one of the key problems highlighted by [8]: artificial neural networks based on backpropagation require buffering past presynaptic activities and synchronizing them with the corresponding error signals propagated sequentially from higher layers.
@@ 2 Error Diffusion Learning Algorithm (EDLA)
The Error Diffusion Learning Algorithm (EDLA) is a novel approach to neural network learning initially proposed by Kaneko [9]. Although the original model is available only in Japanese, this section provides a comprehensive reinterpretation of EDLA.
@@ 2.2 Learning Rule and Mathematical Formulation
$z_{j}^{p,(1)}=g(a_{j}^{p,(1)})\quad\text{and}\quad z_{j}^{n,(1)}=g(a_{j}^{n,(1)}),$ (3)
where $g(\cdot)$ is the activation function. In subsequent hidden layers (l=2,,L)(l=2,...,L) where $L$ refers to the final hidden layer, the neuron outputs are given by
@@
To reflect these constraints at the start of training, weights are initialized by sampling uniformly from $[0,1]$ for synapses connecting neurons of the same type and from [1,0][-1,0] for those connecting neurons of different types.
The objective of training an EDLA network is to identify the optimal parameter set $\mathbf{W}$ that minimizes a predefined error function $E(\mathbf{W})$, similar to traditional MLPs. Gradient descent is a widely used optimization method in which parameters are iteratively updated in a direction that locally reduces the value of $E(\mathbf{W})$. In practice, most practitioners use stochastic gradient descent (SGD) [10], which computes gradient estimates efficiently using subsets of training data (mini-batches). The EDLA also employs a method based on SGD.
@@
[…]
@@
where $\eta$ denotes the learning rate. Thus, the global error signal directly determines whether weights from positive or negative neurons are adjusted during training.
@@
When the network output is smaller than the target value ($d>0$), the absolute values of the weights originating from positive neurons increase due to the properties $g^{\prime}(a)>0$ and $z>0$. Such weight adjustments amplify the outputs of positive neurons in subsequent layers while simultaneously suppressing the outputs of negative neurons in those layers. Consequently, these modifications reduce the discrepancy between the output and target value of the network. Conversely, when the network output exceeds the target value ($d<0$), the absolute values of the weights originating from negative neurons increase. These weight adjustments decrease the outputs of positive neurons in subsequent layers while simultaneously increasing the outputs of negative neurons. Thus, regardless of the sign of the global error signal, EDLA systematically reduces the difference between the network output and target values through successive weight updates. Importantly, this characteristic learning behavior of EDLA is maintained as long as $g^{\prime}(a)>0$, $z>0$, and the sign of the derivative of the loss function properly encodes whether the network output is below or above the target value. The same squared error loss function was used in all experiments conducted in this study.
For the EDLA, the activation function is typically chosen as the sigmoid function. It constrains the output of the neuron between 0 and 1 and its derivative between 0 and 0.25. By limiting both the neuron outputs and their derivatives, the sigmoid activation function provides stability to the training dynamics, avoiding large weight updates and facilitating stable convergence.
@@ 2.3 Biological Interpretation
[…]
Moreover, EDLA employs a global error-diffusion mechanism in which the error signal $d$ is broadcast across the network. This diffusion is conceptually analogous to biological neuromodulation and closely parallels reward-prediction-error (RPE) signals observed in biological systems, which guide synaptic plasticity via diffuse neuromodulatory pathways (e.g., dopamine) [12], thereby mirroring how neuromodulators broadly regulate synaptic change.
Unlike traditional Hebbian learning—which is typically unsupervised and driven solely by local activity correlations—EDLA integrates a supervised component through the global error signal $d$, modulating synaptic changes according to network performance relative to a target. Conceptually, EDLA aligns with neo-Hebbian three-factor rules in which synaptic change depends on (i) a presynaptic factor, (ii) a postsynaptic factor, and (iii) a systems-level modulatory factor [13, 14]. In EDLA the error $d$ plays this third role: weight updates depend on presynaptic activity $z^{p,(l-1)}_{i}$, a postsynaptic state $g^{\prime}(a^{p,(l)}_{j})$, and the global error signal $d$.
In canonical neuroscientific three-factor formulations, pre- and postsynaptic coincidence first sets a synapse-specific eligibility trace (or tag), and a later modulatory signal (e.g., dopamine/RPE) converts that trace into an actual weight update. Eligibility traces bridge the temporal gap between neural activity and delayed reward signals, which is particularly important in spiking networks where modulatory signals often arrive with delays relative to the activity that made a synapse “eligible” [14]. By contrast, EDLA operates in an artificial-network regime where the global error signal is computed and broadcast immediately at each training step. Accordingly, EDLA may be viewed as a limiting case of three-factor rules in which eligibility traces have effectively negligible lifetime and modulatory delays are absent; this approximation is reasonable for supervised tasks with immediate error availability but limits applicability in scenarios with delayed modulatory feedback.
@@ 2.4 Limitation
The inherent simplicity and approximation underlying EDLA introduce significant constraints, particularly concerning the network’s capability to handle multiple outputs effectively. In the EDLA framework, the global error signal $d$ is uniformly diffused throughout the network architecture. This global error diffusion inherently directs the optimization process toward minimizing the error associated with a single-output node. Consequently, when dealing with single-output scenarios, the EDLA can efficiently and accurately specialize the network weights for a specific output. However, when an EDLA network contains multiple outputs, each output generates its own global error signal, which propagates throughout the network. Such simultaneous diffusion of distinct error signals leads to conflicting weight adjustments, as the network simultaneously attempts to minimize errors for multiple outputs. These conflicting updates inhibit the network’s ability to extract and generalize meaningful features from the input data, thereby undermining learning efficacy.
@@ 2.5 Extending EDLA to Multiple Outputs
The inherent single-output constraint of the EDLA represents a significant limitation when addressing tasks that require multiple outputs. A simple approach to adapting EDLA to overcome this limitation involves constructing multiple EDLA networks arranged in parallel. This parallelized architecture addresses a task that requires $K$ outputs by employing $K$ independent EDLA networks. Each individual network retains the fundamental EDLA architecture and operates autonomously. Input vector $\mathbf{x}$ is simultaneously provided to all networks, ensuring that they process the same input data. Each network independently computes output $y_{k}$, where $k\in\{1,\ldots,K\}$. In classification tasks, if the output $y_{k}$ is the maximum value among the outputs of the EDLA networks, $k$ is regarded as the overall output label.
@@
The weights of each individual EDLA network are independently optimized using the standard EDLA update rule. Consequently, this parallel architecture retains the capability, simplicity, and core principles inherent to single-output EDLA models, effectively extending EDLA’s capability to handle multi-output tasks. However, this parallelization approach increases computational costs because the total number of EDLA networks scales linearly with the required number of outputs. This trade-off highlights the challenge of maintaining the computational simplicity of single-output EDLA networks while contending with the additional resource demands necessary for multi-output learning.
@@
Employing multiple single-output networks in an EDLA framework may seem computationally inefficient compared with traditional multi-output MLP architectures. Nevertheless, this parallelization method is reminiscent of modular structures, which are organizational principles observed in biological neural systems, particularly in the brain [15]. Mountcastle’s early neurophysiological research [16] has demonstrated that neurons within the primate somatosensory cortex are organized into vertically aligned columns, each dedicated to a particular modality or receptive field. Additionally, neurons in the primary visual cortex are organized into modular structures known as hypercolumns, each specializing in processing specific features of visual stimuli such as orientation [17]. Each hypercolumn functions semi-independently, collectively contributing to overall visual perception by handling distinct components of the visual scene. Similarly, the inferior temporal cortex exhibits clusters of neurons that are selectively responsive to complex visual features including distinct object shapes [18, 19]. This clustered arrangement enables the efficient processing of diverse visual elements while simultaneously facilitating effective integration across features and categories. Another notable example is the barrel cortex of rodents, in which neural clusters (called barrels) are dedicated to independently processing tactile inputs from individual whiskers [20, 21]. Each barrel independently processes sensory information corresponding to its associated whisker, and the combined neural activity across the barrels generates a comprehensive tactile representation. These modular and distributed neural processing strategies enable the brain to handle complex and heterogeneous tasks efficiently while preserving both flexibility and robustness. Although the parallelized architecture of EDLA, which consists of multiple single-output networks dedicated to specific outputs, may be computationally less efficient than traditional multi-output MLPs, it reflects these biological modular architectures.
The increased computational cost of the parallelized EDLA is an important practical consideration. However, the increase may be less problematic for brain systems or neuromorphic computing applications. In such systems, neurons and synapses are implemented as individually operating, massively parallel [22, 23]. As a result, increases in neuron and synapse counts do not necessarily translate into proportional increases in wall-clock computation time. Some areas of the brain discussed above, such as the visual cortex and the barrel cortex, employ modular architectures that process complex sensory inputs in real-time. Indeed, researchers have achieved real-time simulation using massively parallel systems [22]. Therefore, while the parallelized EDLA entails a trade-off between biological plausibility and computational efficiency on standard computer systems, it may be particularly well suited to neuromorphic implementations, brain-inspired computing systems, or other massively parallel platforms. However, the increase in the number of unit circuits for neuromorphic systems and the memory cost for parallel computing remain a concern, as each EDLA network requires its own set of weights and parameters.
2.6 Number of parameters in EDLA
For an EDLA network with $L$ hidden layers, $n_{\mathrm{hid}}$ positive neurons and $n_{\mathrm{hid}}$ negative neurons per layer, input size $n_{\mathrm{in}}$, and $n_{\mathrm{out}}$ outputs (implemented as $n_{\mathrm{out}}$ independent single-output EDLA networks), the total number of parameters, $N_{\mathrm{EDLA}}$, is
$N_{\mathrm{EDLA}}$ $=\big(2n_{\mathrm{in}}(2n_{\mathrm{hid}})+(L-1)(2n_{\mathrm{hid}})^{2}+2n_{\mathrm{hid}}\big)\,n_{\mathrm{out}}$ (17)
$=2n_{\mathrm{out}}\,n_{\mathrm{hid}}\big(2n_{\mathrm{in}}+2(L-1)n_{\mathrm{hid}}+1\big).$ (18)
Here the first term, $2n_{\mathrm{in}}(2n_{\mathrm{hid}})$, corresponds to the weights connecting the input layer to the first hidden layer; the second term, $(L-1)(2n_{\mathrm{hid}})^{2}$, accounts for the weights between hidden layers; and the third term, $2n_{\mathrm{hid}}$, corresponds to the weights from the last hidden layer to a single output (there are $n_{\mathrm{hid}}$ positive and $n_{\mathrm{hid}}$ negative neurons, hence $2n_{\mathrm{hid}}$). Multiplying by $n_{\mathrm{out}}$ yields the total number of parameters across all independent single-output EDLA networks. Bias terms are omitted in this theoretical analysis for simplicity. The exact parameter counts reported in Table 1, which are used for experimental comparison, include bias terms.
In contrast, the parameter count for a standard fully connected MLP with $L$ hidden layers, $m_{\mathrm{hid}}$ neurons per hidden layer, input size $n_{\mathrm{in}}$, and $n_{\mathrm{out}}$ outputs, $N_{\mathrm{MLP}}$, is
$N_{\mathrm{MLP}}$ $=n_{\mathrm{in}}m_{\mathrm{hid}}+(L-1)m_{\mathrm{hid}}^{2}+m_{\mathrm{hid}}n_{\mathrm{out}}$ (19)
$=m_{\mathrm{hid}}\big(n_{\mathrm{in}}+(L-1)m_{\mathrm{hid}}+n_{\mathrm{out}}\big),$ (20)
again omitting bias terms. Unlike EDLA, the parameter count of a traditional MLP does not scale strongly with $n_{\mathrm{out}}$ (it appears only in the final layer term).
To compare parameter efficiency, consider the case $n_{\mathrm{hid}}=m_{\mathrm{hid}}$ and assume the network is sufficiently deep and wide so that hidden-to-hidden weights dominate. When $2(L-1)n_{\mathrm{hid}}\gg 2n_{\mathrm{in}}+1$, the EDLA parameter count can be approximated as
$N_{\mathrm{EDLA}}\approx 4\,n_{\mathrm{out}}\,n_{\mathrm{hid}}^{2}(L-1).$ (21)
If $n_{\mathrm{hid}}=m_{\mathrm{hid}}$ and $(L-1)m_{\mathrm{hid}}\gg n_{\mathrm{in}}+n_{\mathrm{out}}$, then
$N_{\mathrm{MLP}}\approx n_{\mathrm{hid}}^{2}(L-1).$ (22)
Thus, in this limit,
$\frac{N_{\mathrm{EDLA}}}{N_{\mathrm{MLP}}}\approx 4\,n_{\mathrm{out}}.$ (23)
An EDLA implemented as $n_{\mathrm{out}}$ independent single-output networks can therefore require roughly $4\,n_{\mathrm{out}}$ times more parameters than a comparable MLP with the same hidden-layer size and depth. This substantial increase in the parameter count can raise memory usage and computational cost during training and inference, particularly for tasks with many outputs. The computational costs may, however, be partially mitigated by parallel implementations or neuromorphic hardware, as noted in the previous subsection. If a multi-output EDLA architecture were developed that shares computations across outputs in future work, it could substantially reduce the parameter overhead compared with the current parallelized approach.
2.7 Stabilization by RMS Normalization
When the activation function has no upper bound (for example, ReLU), individual pre-activation values and the corresponding outputs can grow without bound. Consequently, the magnitude of weight updates depends directly on the activation of the preceding layer. In particular, large pre-activations under ReLU can produce very large updates and thereby destabilize the learning dynamics. To mitigate this problem, Root Mean Square (RMS) normalization [24] is applied to the pre-activations. Normalization techniques such as RMS, Layer Normalization [25], and batch normalization [26] are commonly applied to pre-activations in artificial neural networks.
The RMS is defined as
$\mathrm{RMS}=\sqrt{\frac{1}{2n_{\mathrm{hid}}}\sum_{j=1}^{n_{\mathrm{hid}}}\left((a^{p,(l)}_{j})^{2}+(a^{n,(l)}_{j})^{2}\right)}$ (24)
where $n_{\mathrm{hid}}$ is the number of neurons in each sublayer (so that $2n_{\mathrm{hid}}$ is the total number of hidden neurons in the full layer). The normalized pre-activation $\tilde{a}^{p,(l)}_{j}$ is then given by
$\tilde{a}^{p,(l)}_{j}=\frac{a^{p,(l)}_{j}}{\mathrm{RMS}}\gamma.$ (25)
An analogous expression holds for $\tilde{a}^{n,(l)}_{j}$. The positive scalar $\gamma$ controls the overall magnitude of the normalized pre-activations. In this study, $\gamma$ is fixed to 1.0 because learning $\gamma$ is outside the scope of this work and because $\gamma$ is here treated as a biologically motivated gain parameter. A fixed $\gamma$ scales the pre-activation vector of each layer without changing its direction. RMS normalization is applied only between hidden layers; it is not applied between the input layer and the first hidden layer, nor between the final hidden layer and the output layer.
Incorporating RMS normalization, the weight update rule for a connection from a positive presynaptic neuron to a positive postsynaptic neuron is derived as follows:
$dw^{pp,(l)}_{ji}$ $=\eta dg^{\prime}(\tilde{a}^{p,(l)}_{j})\frac{\partial\tilde{a}^{p,(l)}_{j}}{\partial a^{p,(l)}_{j}}z^{p,(l-1)}_{i}$ (26)
$=\eta dg^{\prime}(\tilde{a}^{p,(l)}_{j})\frac{1}{\mathrm{RMS}}\left(1-\frac{{a^{p,(l)}_{j}}^{2}}{2n_{\mathrm{hid}}\mathrm{RMS}^{2}}\right)z^{p,(l-1)}_{i}$ (27)
The term $1-\frac{{a^{p,(l)}_{j}}^{2}}{2n_{\mathrm{hid}}\mathrm{RMS}^{2}}$ is positive because $2n_{\mathrm{hid}}\mathrm{RMS}^{2}$ is the sum of all squared pre-activations in the layer, which is necessarily greater than any single squared pre-activation $(a^{p,(l)}_{j})^{2}$. Therefore, the sign of the weight update $dw^{pp,(l)}_{ji}$ remains determined by the sign of the global error signal $d$, preserving the fundamental learning mechanism of the EDLA. The update rules for other weight types are derived in an analogous manner.
RMS normalization can be interpreted in the context of biological information processing, relating to mechanisms like gain control and divisive normalization. Gain control [27, 28] modulates neuronal sensitivity according to the overall activity level of a population, thereby helping to maintain firing rates within a stable dynamic range and avoiding saturation. Divisive normalization [29, 30] divides a neuron’s response by a pooled measure of population activity, preventing any single unit from dominating the representation. RMS normalization performs a related operation by scaling pre-activations in proportion to the population RMS; this reduces the magnitude of excessively large updates and thus stabilizes learning.
RMS normalization is selected over Layer Normalization due to a key property of the EDLA architecture. In EDLA, non-negative activation functions such as the sigmoid or ReLU are frequently employed. Layer Normalization standardizes activations to have zero mean and unit variance, a process that involves re-centering the data. This re-centering inevitably shifts many pre-activation values into the negative range. In particular, when using the ReLU activation function, this would cause a significant portion of neurons to output zero, effectively silencing them and impeding the weight updates. RMS normalization avoids this issue by rescaling activations without the re-centering step.
Note that RMS normalization is applied only in the experiments reported in Section 4.5; it is omitted in the other experiments described in this manuscript.